### v1

# Enhancing the Generalization Capability of Skin Lesion Classification Models with Active Domain Adaptation Methods

Abstract
We propose a method to improve the generalization ability of skin lesion classification models by combining self-supervised learning (SSL), unsupervised domain adaptation (UDA), and active domain adaptation (ADA). The main steps of the approach include selection of a SSL pretrained model on natural image datasets, subsequent SSL retraining on all available skin lesion datasets, finetuning of the model on source domain data with labels, application of UDA methods on target domain data, and lastly, implementation of ADA methods. The efficacy of the proposed approach is assessed across ten skin lesion datasets of domains, demonstrating its potential for enhancing the performance of skin lesion classification models. This approach holds promise for facilitating the widespread adoption of medical imaging models in clinical settings, thereby amplifying their impact.
1 Introduction
The past decade has witnessed a surge in the application of artificial intelligence (AI) to skin cancer detection. Given the growing incidence of skin cancer within an aging population and the scarcity of dermatology experts, there is a compelling demand for automated AI solutions [60]. Globally it’s estimated that around three billion people lack adequate access to skin disease medical care [15]. Using AI can greatly alleviate the problem. Dermoscopy images are normally used by dermatologists to detect skin cancers. So computer vision (CV) methods have been applied for skin cancer classification and segmentation [54, 38, 37]. Since the start of annual challenges organized by the International Skin Imaging Collaboration (ISIC) in 2016, steady progress has been made in algorithm performance, with deep learning models surpassing human experts on the provided datasets [60, 14].
Despite these advancements in medical imaging applications, model generalization remains a crucial hurdle for broad clinical adoption [35, 69, 14, 26]. For skin cancer detection specifically, Barros et al. and Daneshjou et al. find that models perform worse on dark skin tones than light skin tones, and worse on uncommon disease types [5, 16]. Han et al. observe that the dermatologist-level model’s performance significantly decreases when testing on out-of-distribution data [26]. The differences between skin lesion datasets are pervasive, stemming from various sources such as demographics, skin lesion locations, institutions, acquisition devices, and lighting conditions. Transfer learning (TL) is a technique to transfer a model’s knowledge from one domain to others. Due to the lack of large labeled datasets, a common approach is to initially train models on natural image datasets, such as imagenet [17, 47, 49]. Subsequently, the model is fine-tuned on medical image datasets. There are various factors affecting a model’s transferability, including transfer techniques, training data size, domains’ similarity, and model architecture [46]. Here we focus on the transfer techniques, since some of the factors are out of our control in actual use cases, e.g. we have to work with a given trained model on a given dataset.
One transfer learning technique is using self-supervised learning (SSL). Supervised learning (SL) encourages models to learn task-specific features, potentially hindering their ability to generalize to unseen data. SSL, on the other hand, enables models to learn general features that are more robust across diverse datasets [2]. Azizi et al. propose a robust and data-efficient generalization framework for medical imaging applications by doing SSL training on in-domain medical data prior to finetuning. This approach outperforms the SL models on both in-domain and out-of-domain data. Haggerty et al. confirm the efficacy of this approach on skin lesion datasets, demonstrating that SSL retraining on skin lesion data enhances the performance of a model pretrained on imagenet data with SL or SSL methods [25]. Another TL approach is domain adaptation (DA). DA methods aim to align the source domain (data on which the model is trained) and the target domain (data the model hasn’t encountered). These are few works on applying unsupervised domain adaptation (UDA) methods for skin lesion classification [65, 9, 24].
In this study, we aim to combine the SSL and DA for better transfer learning. To the best of our knowledge, we are the first to apply this approach for skin lesion classification. More specifically, we use the state-of-the-art (SOTA) SSL method, DINO, which is different from the Barlow Twins SSL method used in [2]. Besides using the UDA method, we also apply active domain adaptation (ADA) [45] methods. This is motivated by the actual clinical setting where the model can get feedback from human experts to iteratively improve performance [63]. There is little work on combining the SSL and DA methods. Zhao et al. combines SSL and AL without domain adaptation for skin lesion segmentation application [71]. Xu et al. combines SSL and UDA without ADA, and they use the SSL method of predicting rotations [67]. We demonstrate the effectiveness of the approach in the skin lesion classification application. This approach can easily integrate into the clinical workflow and iteratively improve the model performance by incorporating information from human experts. The approach consists of the four main steps with the initial preparation step of selecting a SSL pretrained model on imagenet. 1. Apply SSL retraining on all available skin lesion datasets. 2. Finetuning the model on the source domain data with labels. 3. Apply unsupervised domain adaptation on the target domain data. 4. Apply active domain adaptation as needed. This step requires input from human experts. They can decide how many labels to provide.
Our main contributions include:
  • We propose a workflow to iteratively improve the generalization capacity of the skin lesion classification model across domains by combining SSL training, UDA and ADA methods.
  • We do an ablation study to find the performance contribution of each step of the workflow, the SSL training, UDA and ADA.
  • Through extensive experiments, we show that our approach achieves SOTA generalization performance on ten skin lesion target domains’ data.
2 Related work
2.1 Self-supervised learning
SSL is a technique to train models without labels. It can be used to create foundational models which can be finetuned for different downstream tasks. The success of SSL is initially observed in natural language processing (NLP) [18, 6], and later extended to CV [19, 20, 23, 39, 50]. Recently there has been a rapid proliferation of SSL applications in the medical imaging field [3, 41, 55, 56, 66, 32]. SSL can achieve better performance than SL trained models with much less labels [10]. There are four main types of SSL methods, innate relationships, generative, contrastive and self-prediction. Innate relationships leverage the internal structure of the data, e.g. predicting the rotation angle, or finding the position of shuffled image patches. Generative methods use autoencoders or generative adversarial networks (GANs). They are assessed by the quality of reconstructing the original images. Contrastive methods rely on using data augmentation to create positive pairs of an image. Different images are treated as negative pairs. There are various distance metrics being proposed trying to keep positive pairs close and negative pairs far away in the embedding space. Some examples include SimCLR [10], MoCo [27], BYOL [4], SimSiam [11], DINO [7]. Self-prediction methods mask out portions of the original image and try to reconstruct the original image. The main difference between self-prediction and generative methods is that self-prediction applies augmentation on portions of the original image while generative methods apply to the whole image.
SSL application in medical imaging is rapidly rising, since medical labels are costly and time consuming to get. SSL can speed up the process of model development. When applying SSL for medical imaging, special attention is needed when choosing the appropriate data augmentation methods. In general, Morningstar et al. shows that augmentation diversity contributes more to SSL performance compared to algorithmic changes [64]. However, due to the difference between natural images and medical images, Morita et al. shows that employing simple transformations including the left/right, up/down flipping, rotation and region cropping with a large ratio of the raw image can improve the classification performance by more than 3% compared with more complicated data augmentations [48].
2.2 Unsupervised domain adaptation
UDA methods can be divided into two types, moment matching and adversarial training. Moment matching methods try to match the first or second moments between the hidden activation distributions of the source and target domains [70]. Some methods include deep adaptation networks (DAN) [42], joint adaptation networks (JAN) [43], correlation alignment [58], and central moment discrepancy (CMD) [70]. Adversarial training methods try to learn domain invariant features. It trains a domain classifier to put the source and target domain features in the same space. The most used adversarial training method is domain adversarial neural networks (DANN) [22]. Other methods build upon it such as adversarial discriminative domain adaptation (ADDA) [62], maximum classifier discrepancy (MCD) [52], batch spectral penalization (BSP) [12] and minimum class confusion (MCC) [34].
There are works of applying UDA in medical image analysis focusing on classification [1], object localization, and segmentation [36]. More specifically for skin lesion classification, Chamarthi et al. compare performance of different UDA methods [9]. Wang et al. show applying UDA methods can mitigate model bias against minority groups [65].
2.3 Active domain adaptation
UDA methods’ performance is often worse than their supervised counterpart [59, 13]. Supervised domain adaptation (SDA) requires labels of the target domain. However, for the medical imaging field especially, obtaining large amounts of labels can be cost intensive. Thus how to select the most informative samples to annotate becomes important. This is where active learning (AL) comes in. AL is trying to maximize a model’s performance given a limited amount of labels. Active domain adaption (ADA) combines active learning with domain adaptation. Different active learning methods have different sampling strategies. Broadly they can be separated into a few categories, such as uncertainty [40, 53], diversity [33, 30], representativeness [68]. There are methods trying to combine multiple types of metrics [51, 31, 21]. Zhao et al. applied SSL and AL for skin lesion segmentation, but there is no domain adaptation [71]. There is no prior work on applying ADA methods to skin lesion classification problems.
3 Methodology
Here we propose the combination of SSL and DA methods into one workflow to improve skin lesion classification model’s generalization ability. It consists of the four steps.
Figure 1: Proposed workflow for model generalization. There are 4 steps, SSL retraining on in-domain data of a SSL pretrained model , finetuning the model on source domain, applying UDA, and lastly applying ADA.
The preparation step is selecting a SSL pretrained model on natural image datasets, e.g. imagenet. Here we select the DINO SSL method since it’s shown to produce the SOTA performance on imagenet among many SSL methods, e.g. SimCLR, MoCov2, Barlow Twins, BYOL, SwAV [44, 7]. It uses knowledge distillation with a student and teacher architecture. The loss term minimizes the cross-entropy between the student and teacher learned features. It applies centering and sharpening on the teacher’s output to avoid learning trivial solutions. It doesn’t require large batch size, or negative samples. We choose to use the ResNet50 [28] backbone model architecture since it’s used in other similar studies [9], and we can compare results. The DINO cross-entropy loss term between the teacher and student learned features:
$minH(P_{t}(x),P_{s}(x))=\min{-(P_{t}(x)\log(P_{s}(x)))}$
$P_{t}(x),P_{s}(x)$ are the probability distribution from the output of the teacher and student network respectively.
The first step is to continue DINO training on skin lesion datasets. Since no label is needed, we can train it on all the available skin lesion data. This can help the model generalize to skin lesion data from natural images.
The second step is finetuning the model on skin lesion dataset with labels. Here we choose a very common task, binary classification of melanoma (cancer) vs. nevus (benign). We keep the backbone model frozen and only finetune the linear classification layer, because we want to directly compare the learned feature quality without the confounding factor of a more complex classifier.
The third step is applying UDA to align the source and target domain. The UDA method we choose is DANN [22]. It’s a classic UDA method and has shown good performance among UDA methods applied to skin lesion datasets [9]. DANN is an adversarial training based method. It has three main components, feature extractor, class classifier, and domain classifier. which tries to learn domain invariant features with a gradient reversal layer.
$L=min_{f},max_{d}Lc(G_{y}(G_{f}(x)),y)+Ld(G_{d}(G_{f}(x)),d)$
Lc and Ld are the cross-entropy loss of skin lesion classification and domain classification receptively. y and d are the true skin cancer label and the true domain label. $Gd,Gy,Gf$ are the networks of domain classifier, class classifier and feature extractor respectively. Their parameters are $d,y,f$ respectively.
The fourth step is applying ADA. We choose the ADA method, AADA [57] since it incorporates the DANN method. The sampling criterion is as such:
$S(x)=\frac{1-G_{d}(G_{f}(x))}{G_{d}(G_{f}(x))}H(G_{y}(G_{f}(x)))$
It includes the diversity cue $\frac{1-G_{d}(G_{f}(x))}{G_{d}(G_{f}(x))}$ and uncertainty cue $H(G_{y}(G_{f}(x)))$. The pseduo code of applying AADA method is show below.
Algorithm 1 AADA
Input: labeled source $L_{s}=H$; unlabeled target $U_{t}\in\{HA,HLH,HLP,B,BA,BLH,BLP,M,MA,MLH\}$; labeled target $L_{t}=\emptyset$; budget per round $b=10$.
Model: =$\mathcal{M}=\{G_{f},G_{y},G_{d}\}$; feature extractor $G_{f}$; class predictor $G_{y}$; discriminator $G_{d}$. $G_{f}$ is DINO trained on skin lesion datasets and finetuned on dataset, H.
UDA: Train $\mathcal{M}$ with ($L_{s}$, $U_{t}$) using DANN loss
for $round\leftarrow 1$ to $MaxRound=10$ do
     Compute $S(x)$ $\forall x\in U_{t}$
     Select a set of $b$ images $z$ from $U_{t}$ according to $S(z)$
     Get labels $y_{z}$ from human experts
     $L_{t}\leftarrow L_{t}\cup(z,y_{z})$
     $U_{t}\leftarrow U_{t}\setminus(z,y_{z})$
     Train $\mathcal{M}$ with ($L_{s}\cup L_{t}$, $U_{t}$)
end for
4 Experiment
4.1 Dataset
[35] provide skin lesion datasets of different domains by grouping data from three datasets, HAM10000 [61], BCN20000 [29], and MSK [8], according to biological traits, such as patient age and lesion location. They quantify the intensity of domain shifts with divergence metrics, cosine similarity and JS divergence. The details of the datasets are in Table 1.
Abbreviation Origin Biological factors Melanoma amount Nevus amount Total target size
H HAM $Age<=30$, Loc. = Body (default) 465 (10%) 4234 (90%) 4699
HA HAM $Age>30$, Loc. = Body 25 (4%) 532 (96%) 557
HLH HAM $Age<30$, Loc. = Head/Neck 90 (45%) 121 (55%) 220
HLP HAM $Age>30$, Loc. = Palms/Soles 15 (7%) 203 (93%) 218
B BCN $Age>30$, Loc. = Body (default) 1918 (41%) 2721 (59%) 4639
BA BCN $Age<=30$, Loc. = Body 71 (8%) 808 (92%) 879
BLH BCN $Age>30$, Loc. = Head/Neck 612 (66%) 320 (34%) 932
BLP BCN $Age>30$, Loc. = Palms/Soles 192 (65%) 105 (35%) 297
M MSK $Age>30$, Loc. = Body (default) 565 (31%) 1282 (69%) 1847
MA MSK $Age<=30$, Loc. = Body 37 (8%) 427 (92%) 464
MLH MSK $Age>30$, Loc. = Head/Neck 175 (60%) 117 (40%) 292
Table 1: Skin lesion datasets of different domains. Here lists a total of 11 datasets. Dataset H is the source domain, and other datasets are target domains. It also shows the number of samples in each dataset, the melanoma and nevus ratio, and biological traits.
4.2 Metrics
We choose to use AUPRC metric because it accounts for unbalanced classes datasets well and it emphasizes more on the cancer class. As we can see from Table 1, many of the skin lesion datasets have very high levels of unbalanced classes with cancer class as low as 4%. We don’t use accuracy or AUROC, because they only work well for balanced classes datasets. However, we do need to be mindful of the different AUPRC baselines of different datasets, which are the melanoma ratios.
4.3 Model Training
When retraining the model with DINO SSL on the in-domain data, we use all the skin lesion datasets from all the domains. We use the default data augmentations, with random cropping, 2 bigger view sizes (0.4-1.0 scale) and 4 smaller view sizes (0.05-0.4 scale), rotation, jitter, blur and black and white transformations. Only the bigger view sizes are passed to the teacher model, while all the view sizes are passed to the student model. This is to encourage learning “local-to-global” correspondence. We have tried different data augmentations, and found that the default setting gives the best result. The training process we find working best is that first, we freeze the backbone model and only train the prediction head layers with a learning rate of 1e-3. We train for 30 epochs. Then we train all the layers with a learning rate of 5e-4. We train for 25 epochs at which point the loss plateaus.
We then finetune the model on the skin lesion source domain dataset, H. We freeze the backbone and only train the classifier head which is a single linear layer, because we want to compare the quality of the backbone features.
Model training is done on a RTX 4090 GPU. The batch size is 128. The image size is 224 x 224.
4.4 Ablation Study
We do ablation study to find the benefits of each step in the workflow, specifically 1. SSL training. 2. Applying the UDA DANN method. 3. Applying ADA (random with finetuning vs. AADA).
5 Results and discussions
Figure 2: Performance comparison of SL and SSL training.
First we compare the generalization capability of a SL pretrained model on imagenet vs. a SSL pretrained model on imagenet. We compare their classification performance on target domains with the AUPRC metric. In Figure 2, we see that the DINO pretrained model gives better results on 8 of the 11 domains, except BLH, BA, and HA which have slightly worse performance. When comparing the DINO pretrained model on imagenet vs. DINO retraining on skin lesion datasets, we see that DINO retraining gives better results on 7 of the 11 domains and on other domains the results are similar. So DINO retraining on in-domain data is helpful.
Figure 3: Effect of applying UDA on different target domains.
In Figure 3, for most of the datasets, with and without UDA give a similar result. But with UDA, it mainly improves for the HLP dataset. When compared with the results reported in [9], our results are better than their UDA results on 6 of the 10 target domains, except on HA, BA, BLH, and MLH having slightly worse results. This shows that SSL pretraining is an effective UDA method, so applying UDA method doesn’t help much. Since it doesn’t decrease the performance and may increase for few datasets, we still apply the additional UDA method.
Figure 4: Effect of applying ADA for 10 iterations.
Here we apply 10 iterations of ADA with each iteration adding 10 samples for annotation. The baseline ADA method is using random selection and finetuning. We choose number 10, since it’s a small and practical number for dermatologists to work with. In Figure 4, we see generally the performance increase with the increasing amount of labels, with the exception of HA and HLP datasets. We see by 3 iterations of ADA with total 30 labels, the performance plateaus. Using AADA and random sampling have a similar performance, but AADA’s performance is more stable as iteration increases.
6 Conclusion
In this paper, we propose a workflow of improving skin lesion classification models’ generalization capability by combining SSL, UDA and ADA methods. Our results show the benefits of each step of the workflow. This workflow helps models to perform well in different domains which mimic the actual clinical use cases, where domain shifts commonly exist.
For future studies, different model architecture can be considered, e.g. vision transformers, since they are shown to give good performance on natural image datasets.